# Computing Large 2D/3D Convex Regions of Obstacle-Free Space for Robotic Planning

### v1

@@ VII-B Comparison of Generating Convex Free Polytope
To begin with, we compare the capability of these methods to adapt to different types of seed and obstacle inputs. As the summarized results in Tab. II shows, FIRI demonstrates the highest level of adaptability to various inputs. Subsequently, based on the reported adaptability, we conduct extensive benchmarks to validate the superior performance of FIRI in terms of managibility, efficiency, and quality.
VII-B1 managibility
@@
When convex polytope is applied in robot navigation, managibility is crucial. For instance, there are situations where we require the convex hull to contain the line segment of the path generated by the frontend [47], or during whole body planning, we demand the convex hull to encompass the robot [48]. Thus we compare the managibility of various methods by using point, line, and convex polytope as the seed input. And the obstacles input are characterized by points.
@@
We conduct benchmark in a complex environment of $50\times 50~{}m$ size (with a height of $10m$ for the 3-D case), where we generate random obstacles by using Perlin noise [49]. For each test, we randomly generate a collision-free seed in the environment as input. The boundary of each convex hull generation algorithm is constrained to be a square with side length $6~{}m$ centered on the seed’s center and parallel to the coordinate axes. And the obstacle input is the points within the boundary of the square in the map. When the seed input is a convex polytope, we set it as a rectangular for the 3-D case, or a rectangle for the 2-D case. We generate the experimental environments with different obstacle densities.
To be fair, we perform several adjustments in different seed cases based on Tab. II. When the seed is a line, for Galaxy and IRIS, which can only use one point as seed input, we use both the endpoints and midpoints of line as seed inputs for them to compute $3$ convex polyhedra. We then select the convex hull with the highest degree of line containment as the result. When the seed is a convex polytope, FIRI can directly take the polyhedra as input. IRIS and Galaxy take the each vertices and center point of the seed as inputs to compute multiple corresponding convex hulls. For RILS, which is adapted to use a line as input, we compute multiple convex hulls using the seed’s edges and diagonals as inputs. For each of the above three methods, we choose the convex hull that maximizes the inclusion of the robot as the respective result.
We calculate the success rate of each method in generating convex polytope that completely contains the seed input, as shown in Tab. III. Due to greedy approach of IRIS in seeking the largest possible volume of convex hull, it performs poorly in terms of managibility, to the extent that it cannot guarantee to contain the seed even when the seed is a point. While RILS can ensure to contains the line seed, it is not sufficient to guarantee that the generated convex hull contain the seed when the seed is represented as a convex hull. As for Galaxy, its heuristic cut method does not ensure that the seed is included either, when the seed is a line or a convex polytope. In contrast, benefiting from Restrictive Inflation, FIRI ensures managibility for all three types of seed inputs across various obstacle densities.
VII-B2 Efficiency and Quality
We construct experiments to compare the efficiency and quality in the aforementioned random environments. The obstacle input and boundary settings remain the same as in Sec. VII-B1. We record the computation time for each algorithm when seeded with a point as input. If IRIS generates a convex polytope during the iteration process that does not include the seed, we force IRIS to terminate prematurely and return the polytope of the previous iteration that includes the seed as the result. Since both RILS and FIRI guarantee managibility for line seeds, we document their respective time overhead when the seed input is a line as well. In addition, FIRI maintains its managibility when only a single iteration is performed, thus we also record the results of a single iteration of FIRI, denoted as FIRI(SI).
@@ VII-B2 Efficiency and Quality
TABLE IV: Comparison of Computation Time between Different Methods for Generating Convex Free Polytope
@@
To provide a fair and intuitive representation of the size of convex polytopes generated by each algorithm, we take FIRI, which aims to maximize the convex hull volume, as the baseline. Specifically, we report the ratio of the volume of the convex hull obtained by each algorithm to the volume of the one obtained by FIRI, as demonstrated in Fig. 7. The efficiency results are presented in Tab. IV, where, for clarity, we distinguish between iterative and non-iterative methods using dashed lines in the case of point seed. In addition, since the number of input obstacles is the primary factor influencing the computational efficiency, we showcase the the number of input obstacles in different density environments in Tab. V to provide readers with an insight into the computational efficiency of each algorithm. As the results Illustrate, both IRIS and FIRI iteratively compute larger convex polytopes by continuously inflating the MVIE, resulting in similar sizes. However, in IRIS, the SDP-based MVIE solving method consumes significant computation time [1], whereas FIRI achieves significant efficiency improvements in MVIE calculations as shown in Sec. VII-A, leading to remarkably more computationally efficient compared to IRIS. Moreover, FIRI achieves a computational time that is within three times the time of non-iterative RILS which even does not require MVIE computations. In contrast, the other three non-iterative methods yield smaller convex polytopes, but achieve better computational efficiency than IRIS in Tab. IV. Galaxy, however, is even less efficient than iterative FIRI due to the fact that the Quickhull [24] algorithm involved degrades in the scene to which Galaxy corresponds, which is particularly noticeable in the 3D case. As for FIRI(SI), similar to RILS, it directly updates the convex hull and can generate polytope of similar size to RILS, but with higher computational efficiency. Moreover, benefiting from the managibility brought about by the Restrictive Inflation, if the objective is to obtain a free convex polytope that encompasses the seed as fast as possible, without pursuing maximum volume, we believe that FIRI(SI) is a suitable choice.
TABLE V: Input Obstacles Number of Different Scenarios
@@
Figure 7: Comparison of the sizes of the free convex polytopes generated by different methods. The dashed line with a value of 1 indicates the result obtained by IRIS, which we take as the baseline.
In order to demonstrate the significance of convex hull quality and its impact on trajectory planning, we conduct an experiment on trajectory planning based on different convex hull generation methods in the random environment as shown in Fig. 8(a). We first generate a collision-free path connecting the start and end points using RRT* [50]. The obtained path can be viewed as a set of connected line segments, based on whcih we generate convex hulls subsequently. As indicated in Tab. III, only FIRI and RILS exhibit managibility over line seed. Thus we build corridors for comparison based on these two methods, which correspond to iterative and non-iterative strategies of whether or not to pursue the maximum volume, respectively. The process of generating a safe flight corridor (SFC) is as follows: Sequentially traverse each line segment on the path. If the line segment is already contained within the previously generated polytopes, proceed to the next segment. Otherwise, we use this line as a seed to generate a new convex polytope. Due to managibility, the generated set of convex polytopes must have an intersection between two neighboring pairs, and this set of convex hull is the result SFC. Based on the generated SFC, we adopt GPOPS-II [51] to obtain the optimal trajectory constrained within the corridor. This collocation-based method transcribe the trajectory optimization problem into a constrained Nonlinear Programming via Gauss pseudospectral method, which is then solved by the well-established NLP solver SNOPT [52]. In GPOPS-II, each trajectory phase is confined within one polytope, and we set the feasibility constraints for velocity and acceleration as $3m/s$ and $6m/s^{2}$, and set the time weight as $20$. As depicted in Fig. 8(b), RILS generates narrow convex polytopes in the area marked by the black dashed circles, resulting in limited maneuvering space for trajectory optimization. Consequently, the trajectory constrained in the SFC generated by RILS exhibits a conservative behavior in the marked aera. In contrast, FIRI, due to its pursuit of maximizing convex hulls, is capable of generating larger corridors, providing greater spatial flexibility in trajectory planning.
@@
(b) The speed profiles of the trajectories, colored by the speed magnitude.
@@
(c) The velocity and acceleration magnitude for the trajectories constrained in the corridors generated by FIRI and RILS respectively.
@@
Figure 8: The comparison between the SFC established based on FIRI and RILS in a complex environment and the optimal trajectories constrained within the generated SFC.
VIII Acknowledgment

### v2

@@ VII-A Comparison of Generating Free Convex Polytope
To begin with, we compare the capability of these methods to adapt to different types of seed and obstacle inputs. As evidenced by the results in Tab. II, FIRI demonstrates the highest level of adaptability to various inputs. Subsequently, based on the reported adaptability, we conduct extensive benchmarks to validate the superior performance of FIRI.
VII-A1 Manageability
@@
As analyzed in Sec. I, manageability is crucial in many applications. For instance, there are situations where we require the convex hull to contain the line segment of the path generated by the frontend [10], or during whole body planning, we demand the convex hull to encompass the robot [11]. Thus we compare the manageability of various methods by using point, line, and convex polytope as the seed. Although the type of obstacles has no effect on manageability, for fairness, the obstacles input are characterized by points.
@@
We conduct benchmark in a complex environment of $50\times 50~{}m$ size (with a height of $10m$ for the 3-D case), where we generate random obstacles by using Perlin noise [55]. For each test, we randomly generate a collision-free seed in the environment as input. The boundary of each convex hull generation algorithm is constrained to be a square (cube for 3-D) with side length $6~{}m$ centered on the seed’s center and parallel to the coordinate axis. And the obstacle input is the points within the boundary of the square in the map. When the seed input is a line, we set its length to $1.5~{}m$. When the seed input is a convex polytope, we set it as a $1.5\times 0.75\times 0.25~{}m$ rectangular for the 3-D case, or a $1.5\times 0.75~{}m$ rectangle for the 2-D case. We generate the experimental environments with different obstacle densities, whose numbers of input obstacles are shown in Tab. IV. For each density, we create different $10$ environments. Furthermore, for each environment, we conduct $500$ random trials for each type of seed input.
@@
TABLE II: Comparison of Adaptability of Different Methods for Generating Free Convex Polytope
Method Adaptability Seed Type Obstacles Type
  Point Line Polytope Point Polytope
FIRI
IRIS[1]
Galaxy[9]
RILS[2]
TABLE III: Comparison of Computation Time of Different Methods for Generating Free Convex Polytope
@@
To be fair, we perform several adjustments in different seed cases based on Tab. II. When the seed is a line, for Galaxy and IRIS, which can only use one point as the seed input, we use both the endpoints and midpoints of the line as seed inputs for them to compute $3$ convex polytopes. We then select the convex hull with the highest degree of line containment as the result. When the seed is a convex polytope, FIRI can directly take the polytope as input. IRIS and Galaxy take each vertex and the center point of the seed as inputs to compute multiple corresponding convex hulls. For RILS, which is adapted to use a line as input, we compute multiple convex hulls using the seed’s diagonals as inputs. For each of the above three methods, we choose the convex hull that maximizes the inclusion of the seed as the respective result.
TABLE IV: Input Obstacles Number of Different Scenarios
@@
Figure 6: Comparison of the sizes of the free convex polytopes generated by different methods. The dashed line with a value of 1 indicates the result obtained by FIRI, which we take as the baseline.
We calculate the success rate of each method in generating a convex polytope that fully contains the seed input, as shown in Tab. I. In addition, for a more intuitive presentation, we present an example of the results of each method in an application of whole-body planning that has a need for manageability. As shown in Fig. 11(b), we use a rectangular robot as the seed, and the details of the application will be described in Sec. VIII-A1. Due to greedy approach of IRIS in seeking the largest possible volume of the convex hull, it performs poorly in terms of manageability, to the extent that it cannot guarantee to contain the seed even when the seed is just a single point. While RILS can ensure to contain the line seed, it is not sufficient to guarantee that the generated convex hull contains the seed when the seed is represented as a convex hull. As for Galaxy, its heuristic cut method does not ensure that the seed is included either, when the seed is a line or a convex polytope. In contrast, benefiting from RsI, FIRI ensures manageability for all three types of seed inputs across various obstacle densities.
VII-A2 Efficiency and Quality
We conduct experiments to compare efficiency and quality in the aforementioned random environments. The obstacle input and boundary settings remain the same as in Sec. VII-A1. We record the computation time for each algorithm when seeded with a point as input. If IRIS generates a convex polytope during the iteration process that does not include the seed, we force IRIS to terminate prematurely and return the polytope from the previous iteration that includes the seed as the result. Since both RILS and FIRI guarantee manageability for line seeds, we document their respective time overhead when the seed input is a line as well. In addition, FIRI maintains its manageability when only a single iteration is performed, we also record the outcomes of a single iteration of FIRI, denoted as FIRI(SI).
To provide an intuitive comparison of the size of convex polytopes generated by each algorithm, we take the proposed FIRI, which aims to maximize the convex hull volume, as the baseline. Specifically, we report the ratio of the volume of the convex hull obtained by each algorithm to the volume of the one obtained by FIRI, as demonstrated in Fig. 6. The efficiency results are presented in Tab. III, where, for clarity, we distinguish between iterative and non-iterative methods using dashed lines in the case of point seed. In addition, since the number of input obstacles is the primary factor influencing the computational efficiency, we showcase the the quantity of input obstacles in different density environments in Tab. IV to provide readers with an insight into the computational efficiency of each algorithm. As the results illustrated in Fig. 6 and Tab. III, both IRIS and FIRI iteratively compute larger convex polytopes by continuously inflating the MVIE, resulting in similar sizes. However, in IRIS, the SDP-based MVIE solving method consumes significant computation time [1], whereas FIRI achieves significant efficiency improvements in MVIE calculations as shown in Sec. VII-C, leading to remarkably higher efficiency compared to IRIS. Moreover, FIRI achieves a computational time that is within three times the time of non-iterative RILS that does not involve MVIE computations. The other three non-iterative methods yield smaller convex polytopes, but achieve better computational efficiency than IRIS in Tab. III. Galaxy, however, is even less efficient than FIRI due to the fact that the Quickhull [30] algorithm involved degrades in the scene to which Galaxy corresponds, which is particularly noticeable in the 3D case. As for FIRI(SI), similar to RILS, it directly updates the convex hull and can generate polytope of comparable size to RILS, but with higher computational efficiency. Moreover, benefiting from the manageability brought about by the RsI, if the objective is to obtain a free convex polytope that includes the seed as fast as possible, without pursuing maximum volume, we believe that FIRI(SI) is a more suitable choice.
Figure 7: The benchmark on computation time with the number of obstacle vertices for FIRI and IRIS in different environmental densities. The top represents the 2-D case and the bottom represents the 3-D case. FIRI has orders of magnitude advantage in computational efficiency over IRIS.
Apart from using point representation for obstacles as used above, employing convex polytopes is also a commonly used method for representing the environment. Among the the above methods shown in Tab. II,, only FIRI and IRIS support obstacle inputs in the form of convex hulls, thus we further compare these two algorithms. For each test we generate a certain number of randomly distributed obstacles in a closed space and select the center of the space as the seed input. This closed space serves as the boundary input to each method. We conducted benchmarks by setting the number $N_{v}$ of the vertices of a single convex obstacle to vary from $3$ to $10^{3}$. Compared to the case where obstacles are represented by points, the relative volumes of obstacles that are convex polytopes are larger, thus we vary the number of obstacles to vary from $10$ to $10^{3}$, which we denote as spare, medium and dense in the results shown in Fig. 7. For each the number of the vertices of each different number of obstacles, we perform $500$ tests. The increasing number of obstacles, the number of halfspaces generated for convex hull construction tends to stabilize, thus the computational efficiency advantage provided by the porposed MVIE solution gradually diminishes. However, benefiting from the efficient SDMN we propose for halfspace computation, FIRI can still maintain a significant advantage over IRIS, even in scenarios with dense obstacles, with a reduction of over 95% in computational requirements.
@@ VII-A2 Efficiency and Quality
(b) The speed profiles of the trajectories, colored by the speed magnitude. FIRI generates larger convex polytopes providing more spatial freedom for trajectory optimization.
@@
(c) The velocity and acceleration magnitude for the trajectories constrained in the corridors generated by FIRI and RILS respectively. Greater spatial freedom provided by FIRI leads to more aggressive trajectory performance.
@@
Figure 8: The comparison between the corridor established based on FIRI and RILS in a complex environment and the optimal trajectories constrained within the generated corridor.
VII-A3 Case Study of Quality
In order to demonstrate the significance of convex hull quality and its impact on trajectory planning, we conduct an experiment on trajectory planning based on different convex hull generation methods in the random environment as shown in Fig. 8(a). We first generate a collision-free path connecting the start and end points using RRT* [56]. The obtained path can be viewed as a set of connected line segments, based on which we generate convex hulls subsequently. As indicated in Tab. I, only FIRI and RILS exhibit manageability over line seed. Thus we build corridors for comparison based on these two methods, which correspond to iterative and non-iterative strategies, regarding the pursuit of maximum volume, respectively. The process of generating a safe flight corridor is as follows: We sequentially traverse each line segment on the path. If the line segment is already included in the previously generated polytopes, we move on to the next segment. Otherwise, we use this line as a seed to generate a new convex polytope. Due to manageability, the generated set of convex polytopes must have an intersection between two neighboring pairs, forming a corridor as the result. Based on the generated corridor, we adopt GPOPS-II [57] to obtain the optimal trajectory constrained within the corridor. This collocation-based method transcribes the trajectory optimization problem into a constrained Nonlinear Programming using the Gauss pseudospectral method, which is then solved by the well-established NLP solver SNOPT [58]. In GPOPS-II, each trajectory phase is confined within one polytope, and we set the feasibility constraints for velocity and acceleration as $3m/s$ and $6m/s^{2}$, and set the time weight as $20$. As depicted in Fig. 8(b), RILS generates narrow convex polytopes in the area marked by the black dashed circles, resulting in limited maneuvering space for trajectory optimization. Consequently, the trajectory constrained in the corridor generated by RILS exhibits a conservative behavior in the marked area . In contrast, FIRI, due to its pursuit of maximizing convex hulls, is capable of generating larger corridors, providing greater spatial flexibility in trajectory planning.
VII-B Comparison of Solving Strictly Convex Small Dimensional Minimum-Norm with Massive Constraints
For the minimum-norm (16), we compare SDMN which is proposed in Sec. IV with several cutting-edge general-purpose QP solvers: the parametric active-set algorithm qpOASES [14], the operator-splitting based first-order approximation algorithm OSQP [15] and the interior-point method-based quadratic programming approximation algorithm HPIPM [16]. We use the open source code  [^1] https://github.com/coin-or/qpOASES [^2] https://github.com/osqp/osqp [^3] https://github.com/giaf/hpipm of these solvers to obtain the solution of (16) respectively. For qpOASES and HPIPM, we use the default parameters. However, we find that the default parameters of OSQP are insufficient when the constraint size becomes large. Therefore, we adjust the relative convergence tolerance to $10^{-12}$ and the maximal number of iterations to $10^{6}$, denoted as OSQP, while the implementation with the default parameters are denoted as OSQP default. As for the proposed SDMN, we do not require any additional parameter settings.
We compare the solving time of each method, as well as their average precision which is defined as
$\psi_{N}=\mathcal{L}_{\psi}({Ey^{*}-f}),$ (61)
where $y^{*}$ is the corresponding solution of each method, $E$ and $f$ are the parameters used to define the constraints of the minimum-norm (16), and $\mathcal{L}_{\psi}(\cdot)$ denotes a function that takes the absolute value of the largest element of the input vector. The precision $\psi_{N}$ indicates the degree of proximity of the solution of each method to the most active constraints of the input minimum-norm. The performance is reported in Fig. 9 and Tab. V. The results demonstrate that SDMN exhibits remarkably high computational efficiency, surpassing other methods by orders of magnitude. Additionally, SDMN provides analytical solutions, achieving high accuracy without the need for additional parameters.
Figure 9: Computation time $t_{comp.}$ of different methods for solving strictly convex small dimensional QP under different constraint numbers. The proposed SDMN outperforms other methods by orders of magnitudes.
TABLE V: Comparison of Precisions $\psi_{N}$ between Different Methods for Solving Small Dimensional Minimum-Norm
Scenario Precisions $\psi_{N}$
qpOASES[14] OSQP[15] default OSQP[15] HPIPM[16] SDMN (Proposed)
2-D 1.28e-15 5.05e-2 6.61e-11 7.79e-06 2.78e-17
3-D 2.91e-15 2.43e-2 2.95e-11 4.41e-06 3.55e-17
VII-C Comparison of Solving MVIE
For MVIE (11), we propose two algorithms, namely the SOCP-reformulation algorithm and the randomized algorithm specialized for 2-D case, which are presented in Sec. V and VI, respectively. The randomized algorithm yields an analytical solution, here we abbreviate it as RAN. Then to distinguish from the methods to be compared, we denote the proposed SOCP-reformulation algorithm which is solved by affine scaling method as CAS. We compare them with three methods: 1) The optimization method based on SDP formulation of MVIE in IRIS [1]. 2) The example  [^4] https://docs.mosek.com/latest/cxxfusion/examples-list.html#doc-example-file-lownerjohn-ellipsoid-cc of the cutting-edge solver Mosek [23] to computes the Lowner-John inner ellipsoidal approximations of a polytope. 3) A strategy of solving the SOCP form (35) of the MVIE by Mosek. As the example in Mosek formulate MVIE into a mixed conic quadratic and semidefinite problem, we denote it as Mosek SDP and denote another strategy using Mosek as Mosek SOCP. We use the default parameters for IRIS and Mosek.
Figure 10: Computation time $t_{comp.}$ of different methods for solving MVIE under different constraint numbers. The proposed CAS outperforms other methods by orders of magnitudes. Additionally, the proposed analytical method RAN further enhances efficiency for the 2D case.
TABLE VI: Comparison of Precision $\psi_{\mathcal{E}}$ between Different Methods for Solving MVIE
Scenario Precisions $\psi_{\mathcal{E}}$
IRIS[1] Mosek[23] SDP Mosek[23] SOCP CAS (Proposed) RAN (Proposed)
2-D 8.56e-8 6.47e-9 4.87e-12 1.59e-8 4.41e-16
3-D 1.54e-8 1.56e-8 4.05e-12 2.04e-8 /
where ${A_{\mathcal{E}}}^{*}$, ${D_{\mathcal{E}}}^{*}$ and ${b_{\mathcal{E}}}^{*}$ are the coefficients of the maximum inscribed ellipsoid solved by each method, $A_{\mathcal{P}}$ and $b_{\mathcal{P}}$ are the parameters that define the halfspaces of the input polytope $\mathcal{P}$ defined in (10), and $\mathcal{L}_{\psi}(\cdot)$ is the function defined in (61). This precision $\psi_{\mathcal{E}}$ represents the proximity of the ellipsoids obtained by different methods to the most active halfspace constraint, which are constructed in the form of second-order cone constraint as (27b). The results are summarized in Fig. 10 and Tab. VI. By comparing the performance of Mosek SDP and Mosek SOCP, we can get the conclusion that transforming the commonly used SDP formulation to the SOCP formulation, which is presented in Sec. V-A, leads to a significant improvement in computational efficiency without sacrificing precision. Additionally, in the low-dimension massive-constraint case faced in this paper, the use of the affine scaling method avoids the requirement of solving a large-scale system of linear equations at each iteration, compared to the primal-dual interior point method used in Mosek. Therefore as the results demonstrate, CAS is capable of solving MVIE in less time while maintaining a comparable precision. Moreover, for the 2-D case, our proposed linear-time complexity algorithm RAN further enhances the computational efficiency by orders of magnitude and compute analytical results.
VIII Real-world Application
(a) Left: Illustration of the generation of a dense corridor based on the yellow rough trajectory obtained from the front-end. Right: We optimize a smooth safe whole-body trajectory by constraining the trajectory in the generated corridor in the back-end.
(b) Left: For the two specific cases in (a), we demonstrate the convex polytopes generated by the different methods. The results show that only FIRI is able to satisfy the whole-body planning requirement of generating polytopes that fully contain the robot. Right: Snapshots of the robot’s navigation result.
Figure 11: A real-world application of a differential driven robot tasked with traversing a maze, utilizing FIRI to abstract the feasible region.
Figure 12: (a)-(e): Illustration of the corridors generated by FIRI during the traversal of the robot from $G_{1}$ through $G_{2}$-$G_{5}$ and finally back to $G_{1}$ in a cluttered environment. (f): Snapshots of the robot’s navigation result.
To validate the performance of FIRI in practical applications, we designed two real-world applications: a global trajectory planning for a vehicle robot with nonholonomic constraints in 2-D environment and a local planning for a quadrotor in 3-D environment.
VIII-A Dense Corridor for 2-D Whole-body Planning
The trajectory planning for large-volume vehicle in narrow environments requests the entire vehicle to be constrained within the safe region, which puts a demand on manageability.
We adopt a differential driven AgileX SCOUT MINI  [^5] https://global.agilex.ai/chassis/11 as the physical platform. The platform is equipped with an Intel NUC  [^6] https://simplynuc.com/product/nuc11phki7c-full/ with an Intel Core i7-1165G7 as onboard processor and a LiDAR sensor  [^7] https://ouster.com/products/hardware/os1-lidar-sensor integrated with an IMU for perception. We use FAST-LIO2 [59] for both pre-mapping the environment and real-time localization during execution of the planned trajectory. That is, the obstacles are represented by points.
Building on our previous work [11], we use hybrid A* as the front-end to obtain an initial feasible trajectory, and then finely sample on the trajectory. The robot shape corresponding to each sampled state is used as a seed to generate a convex polytope by FIRI, forming a dense safety corridor, as shown in Fig. 11. In the back-end, we utilize polynomial representation for the trajectory and constrain that the vehicle shape corresponding to each sampled constraint state on the trajectory is fully contained within the corresponding convex polytope of the safety corridor. Considering localization errors and control errors, for safety purposes, we inflate the robot’s shape during planning to generate a safer corridor and trajectory.
VIII-A1 Narrow Maze
We task the robot with traversing through a narrow maze. The result, as depicted in Fig. 11(a), demonstrates that FIRI, based on the yellow front-end trajectory, generates high-quality convex polytopes, providing sufficient spatial degrees of freedom for trajectory optimization in the back-end. Additionally, as shown in Fig. 11(b), to intuitively demonstrate the manageability of our experiment, we select two particularly narrow cases within the maze environment, illustrating the results of the convex polytopes generated by various methods [1, 9, 2] compared in Sec. VII-A. It can be observed that only FIRI is capable of fully enclosing the robot. A more detailed demonstration of the generation from each seed and its corresponding convex polytope along the entire trajectory is presented in the accompanying video. In this experiment, we generate a dense corridor consisting of $550$ convex polytopes. For each convex polytope, the average number of input obstacle points is $1307.1$, with an average processing time of $0.249~{}ms$. The total time taken for generating the entire corridor is $136.74~{}ms$.
VIII-A2 Cluttered Environment
The results are shown in Fig 12. We require the robot to navigate through a cluttered environment filled with irregular obstacles, starting from $G_{1}$ and sequentially reaching $G_{2}$-$G_{5}$, and finally returning to $G_{1}$. Since the focus of this paper is on the generation of free convex polytope, only the front-end trajectory that guides the generation of convex polytope and the generated convex polytopes are demonstrated. We refer the readers to the accompanying video for the detailed navigation process. In this experiment, as shown in Fig 12, we generate $5$ dense corridors, with an average of $91.8$ convex polytopes in each corridor. For each convex polytope, the average number of input obstacle points was $715.3$, with an average processing time of $0.131~{}ms$. The average time taken to generate a corridor is $12.02~{}ms$.
VIII-B Sparse Corridor for 3-D Local Replanning
Our experimental platform is a customized quadrotor equipped with an NVIDIA Orin NX [^8] https://www.nvidia.com/en-us/autonomous-machines as the onboard processor and a Livox MID-360999https://www.livoxtech.com/mid-360 LiDAR for perception. For the quadrotor, due to its small size, we model it as a point mass, similar to many other works [2, 8]. We employ FAST-LIO2 [59] for online state estimation and utilize occupancy grid map to filter sensor noise, which is easy to inflate for safety margins. For trajectory planning, the path generation and corridor generation are the same as Sec. VII-A3. Then building upon our previous work [3], we utilize piecewise polynomials to represent the trajectory. Each piece of the trajectory is constrained within the corresponding polytope to ensure safety.
As shown in Fig. 13, the experiment is conducted in a dense forest, and the quadrotor is required to navigate through a series of waypoints. Since the environment is unknown, we make the quadrotor perform high-frequency replan (20Hz) according to the real-time perception. The maximum velocity of the UAV during flight is up to 4.5 m/s. Each replan involves 3-7 convex polytopes, the average generation time for each polytope is $2.76~{}ms$, and the average number of input obstacle points for FIRI is $8219.6$. We refer readers to the accompanying video for more information.
IX Discussion and Conclusion
Our algorithm FIRI aims to generate the largest possible convex polytopes. However, when it comes to practical applications such as trajectory planning, we cannot guarantee that the inflation direction of the polytope will always be favorable for trajectory optimization. For example, when the robot has a high velocity, if the inflation direction is perpendicular to the velocity direction, it may lead to the generated convex polytope that are not conducive to the trajectory optimization. Our future research direction will focus on generating convex polytopes that are application-friendly.
In conclusion, we propose a novel obstacle-free convex polytope generation algorithm called FIRI, which achieves high quality, efficiency, and manageability simultaneously. To achieve efficiency, we design targeted methods for the two optimization problems involved in FIRI. Specifically, for 2-D MVIE, we develop a linear-time complexity method, which is proposed for the first time. We perform extensive benchmarks against several convex polytope generation algorithms to confirm the superior performance of FIRI. The comparisons with general-purpose solvers demonstrate the speedup over orders of magnitude of our targeted methods. Two typical applications showcase the practicality of FIRI.
Figure 13: Illustration of the application of FIRI in a local trajectory planning framework for a quadrotor to navigate through a cluttered forest. Top: A snapshots of the quadrotor traveling through the forest. Bottom: The visualization of quadrotor replanning for the moments corresponding to the red boxes. The black transparent polytopes are generated by FIRI.
In CGAL [60], there has been a lack of a dedicated solver for 2-D MVIE. In the future, we will continue our research on 2-D MVIE and work towards implementing a version of rational predicates.
X Acknowledgment